# Tracking the Unstable: Appearance-Guided Motion Modeling for Robust Multi-Object Tracking in UAV-Captured Videos

### v1

@@ Abstract
Multi-object tracking (MOT) aims to track multiple objects while maintaining consistent identities across frames of a given video. In unmanned aerial vehicle (UAV) recorded videos, frequent viewpoint changes and complex UAV-ground relative motion dynamics pose significant challenges, which often lead to unstable affinity measurement and ambiguous association. Existing methods typically model motion and appearance cues separately, overlooking their spatio-temporal interplay and resulting in suboptimal tracking performance. In this work, we propose AMOT, which jointly exploits appearance and motion cues through two key components: an Appearance-Motion Consistency (AMC) matrix and a Motion-aware Track Continuation (MTC) module. Specifically, the AMC matrix computes bi-directional spatial consistency under the guidance of appearance features, enabling more reliable and context-aware identity association. The MTC module complements AMC by reactivating unmatched tracks through appearance-guided predictions that align with Kalman-based predictions, thereby reducing broken trajectories caused by missed detections. Extensive experiments on three UAV benchmarks, including VisDrone2019, UAVDT, and VT-MOT-UAV, demonstrate that our AMOT outperforms current state-of-the-art methods and generalizes well in a plug-and-play and training-free manner. The source codes will be released.
I Introduction
Multi-object tracking (MOT) is a fundamental vision task with widespread applications, such as autonomous driving [1] and unmanned aerial vehicle (UAV) surveillance [2]. A typical pipeline of MOT is to first detect multiple objects, and then assign each detection to existing tracks through data association, ensuring the continuity of track identities over time. Despite progress, robust data association remains challenging, particularly for videos captured by UAV-mounted cameras.
@@ I Introduction
To address the above issues, prevailing methods adopt two main strategies for constructing the cost matrix. (1) Motion-based position prediction: The Kalman Filter [3, 4] is widely used as a motion model to predict current positions of tracks from their historical states. Then, a motion-based cost matrix is constructed based on the positional proximity between detections and tracks. To overcome the limitations of the Kalman Filter’s linear motion predictions in UAV views, some recent works introduce complementary techniques, such as camera motion compensation (CMC) [5, 6] and optical flow [7, 8], to improve the accuracy of predicted track positions. (2) Appearance-based instance-level discrimination: Instance-level re-identification (ReID) embedding, being insensitive to position changes, offers significant advantages under substantial camera motion or object displacement. Several works [9, 5, 10, 6] have constructed appearance-based cost matrices using ReID embedding to facilitate accurate identity assignment. Nevertheless, these two strategies model motion and appearance cues independently to generate separate cost matrices, ignoring the intrinsic relationships between them. Specifically, motion prediction errors caused by sudden object displacements can adversely affect the construction of the motion-based cost matrix, while appearance ambiguities similarly impact the appearance-based cost matrix. Consequently, when these cost matrices produce conflicting association scores, determining a reliable matching decision becomes difficult.
@@
Together with the AMC and MTC modules, we propose a novel multi-object tracker, termed AMOT, which is built on the joint detection and embedding (JDE) architecture [11]. AMOT is specifically designed to enhance the robustness and accuracy of identity assignment under challenging UAV-captured videos. Experiments on multiple UAV benchmarks demonstrate that AMOT achieves superior performance in terms of IDF1 and MOTA, as shown in Fig. 1. The main contributions are summarized as follows:
@@
II Related Works
II-A Motion Modeling
@@
The purpose of motion modeling in MOT is to predict the positions of tracks. Most MOT methods [12, 13, 14] adopt the Kalman Filter for motion modeling, owing to its high computational efficiency for real-time applications. Predicted positions of tracks are compared with current detections via Mahalanobis distance [15] or Intersection over Union (IoU) [16], which are employed as motion-based cost metrics for data association. Some works [4, 17] enhance the Kalman Filter by adopting an observation-centric strategy to refine track estimation, yielding better performance under non-linear motion patterns. Despite their effectiveness, Kalman-based methods exhibit limited capability in scenarios involving substantial camera motion and object displacement. Meanwhile, learning-based motion models [18, 19, 20] have recently gained attention. These models leverage data-driven architectures to learn motion patterns. However, such methods often incur high computational costs and are unsuitable for real-time tracking. In contrast, we introduce a training-free motion modeling strategy that balances robustness and efficiency, specifically tailored for UAV-captured videos.
II-B Appeareance Modeling
@@
Appearance modeling aims to extract discriminative appearance features to re-identify objects. The tracking-by-detection paradigm [21, 22] first detects objects using an off-the-shelf detector, and then employs a ReID network to extract identity embedding for each object. Although this pipeline delivers impressive performance, it suffers from high computational cost due to separate detection and extraction stages. In contrast, the joint detection and embedding (JDE) paradigm [23, 24] integrates object detection and ReID feature extraction into a unified framework, enabling simultaneous object localization and embedding extraction. Furthermore, several JDE-based trackers incorporate global attention [25, 26] and temporal cues [27, 28] to enhance the discriminability of instance-level ReID embedding. The aforementioned methods measure instance-level appearance similarity by computing the cosine distance between ReID embeddings. Differently, we reformulate the appearance similarity measurement as a dense response between the instance embedding and the global ReID feature map. This enables the joint modeling of visual similarity and spatial coherence.
@@
Figure 2: Tracking pipeline of AMOT. Specifically, we introduce the appearance-motion consistency (AMC) matrix $\mathbf{C}_{AMC}$ that integrates it with the appearance similarity matrix $\mathbf{C}_{App}$ and the Intersection-over-Union (IOU) matrix $\mathbf{C}_{IOU}$ to robustly associate high-confidence detections $\mathcal{D}_{h}$ with tracks $\mathcal{T}$ at frame $t-1$. Then, the unmatched tracks $\mathcal{T}_{un}$ are associated with low-confidence detections $\mathcal{D}_{l}$ in the second stage. The remaining unmatched tracks $\mathcal{T}_{second\_un}$ are further potentially reactivated through our proposed motion-aware track continuation (MTC) module. KF means the Kalman Filter.
II-C Data Association
@@
Data association typically involves constructing a cost matrix that quantifies the affinity between current detections and existing tracks based on motion and appearance information. To improve association accuracy, motion and appearance cues are often modeled separately and then integrated into a unified cost matrix [29, 30]. Once the cost matrix is constructed, data association is formulated as an assignment problem and solved using the Hungarian algorithm [31, 32]. Nevertheless, existing approaches commonly fail to consider the inherent interplay between motion and appearance cues. This independent modeling can cause instability in affinity measurement, arising from prediction errors or appearance ambiguities, which ultimately leads to conflicting data associations. To this end, we propose the AMC matrix, which jointly enforces consistency in appearance, motion, and temporal domains, enabling reliable identity assignment.
III Methodology
III-A Preliminaries
@@
where $\mathcal{O}_{det}$ denotes the set of center coordinates for detections, and $N$ is the total number of detections. For each detection, the width and height are obtained from the regression map as $[w_{i},h_{i}]=\textbf{B}^{t}_{(x_{i},y_{i})}$, and the confidence score is given by $c_{i}=\mathbf{H}^{t}_{(x_{i},y_{i})}$. Meanwhile, the ReID embedding $\mathbf{e}_{i}$ for each detection is extracted from $\textbf{E}^{t}$ at the corresponding center coordinate, expressed as $\mathbf{e}_{i}=\mathbf{E}^{t}_{(x_{i},y_{i})}\in\mathbb{R}^{D}$.
@@
In our AMOT, as shown in Fig. 2, we construct an appearance-motion consistency (AMC) matrix to enable robust detection-to-track association. Additionally, we propose a motion-aware track continuation (MTC) module designed to reactivate unmatched tracks without relying on explicit detection-to-track matching.
III-B Appearance-Motion Consistency Matrix
@@
[…]
@@
Subsequently, as shown in Fig. 3, we quantify the appearance-guided spatio-temporal correspondence by measuring the forward and backward spatial distances, which can be defined as:
@@
III-C Motion-aware Track Continuation
@@
III-D Tracking Pipeline
The tracking pipeline of AMOT is illustrated in Fig. 2. The Kalman Filter is used to predict the current positions of tracks. We then adopt a two-stage matching strategy similar to BYTE [3], where detections are divided into high-confidence sets $\mathcal{D}_{h}$ and low-confidence sets $\mathcal{D}_{l}$.
Specifically, in the first stage, we construct three pairwise cost matrices between $\mathcal{D}_{h}$ and the tracks $\mathcal{T}$, including the appearance similarity matrix $\mathbf{C}_{App}$ [11], the Intersection-over-Union (IOU) matrix $\mathbf{C}_{IOU}$, and the proposed AMC matrix $\mathbf{C}_{AMC}$. These matrices are integrated into a unified cost matrix $\mathbf{C}_{uni}$ as follows:
@@ III-D Tracking Pipeline
IV Experiments
IV-A Settings
IV-A1 Datasets and Evaluation Metrics
We evaluated our method on various MOT datasets from UAV perspectives, including VisDrone2019 [33], UAVDT [34], and VT-MOT-UAV [35].
We adopt the widely-used CLEAR metrics [27], including MOTA, IDF1, the number of mostly tracked (MT) objects, and mostly lost (ML) objects, to comprehensively evaluate the tracking performance. MOTA emphasizes the detection quality and is computed based on false positives (FP), false negatives (FN), and identity switches (IDs). IDF1 measures the tracker’s ability to maintain consistent object identities over time, reflecting the accuracy of identity association across frames. Additionally, we report FPS to assess the inference speed of the tracker.
IV-A2 Implementation Details
We adopt DLA-34 [11] as the default backbone network and initialize its parameters with pre-trained weights from the COCO dataset. The input image is resized to 608×1088608$\times$ 1088608 × 1088. The corresponding feature map undergoes a 4$\times$ downsampling, resulting in a resolution of 152×272152$\times$ 272152 × 272, where $H=152$ and $W=272$. We employ the Adam optimizer with an initial learning rate of $7e^{-5}$ to train our model for 30 epochs, reducing the learning rate by a factor of 10 at 20 epochs. For the objective functions, focal loss is employed to supervise the object heatmap, while L1 loss is used to supervise the predicted object width and height. In addition, both cross-entropy loss and triplet loss are applied to guide the learning of the ReID embedding. To further enhance the model’s perception of object instances, we introduce an additional mask branch [36] during training, which is trained using the Dice loss.
Our tracker is implemented using Python 3.7 and PyTorch 1.7.1. All experiments are evaluated on an NVIDIA RTX 3090 GPU and Xeon Platinum 8375C 2.90GHz CPU.
IV-B Benchmark Evaluation
We present the tracking results for multiple UAV datasets. $\uparrow$/$\downarrow$ indicate that higher/lower is better, respectively. The best scores for each metric are marked in bold, and the second-best scores are marked in underline.
IV-B1 Results on VisDrone2019
VisDrone2019 serves as a fundamental benchmark for multi-object tracking in dynamic UAV-captured videos, which involve scale variations and long-range object movements. As shown in Table I, our AMOT exhibits the highest IDF1 of 61.4% and MOTA of 46.0%, outperforming current state-of-the-art methods. Concretely, compared with DroneMOT and MM-Tracker, AMOT achieves absolute improvements in both IDF1 and MOTA, demonstrating superior association accuracy and enhanced overall tracking performance.
TABLE I: Tracking performance comparison with state-of-the-art methods on VisDrone2019 test set.
@@ IV-B1 Results on VisDrone2019
ByteTrack [3] 37.0 35.7 - - 2168
UAVMOT [27] 51.0 36.1 520 574 2775
OC-SORT [4] 50.4 39.6 - - 986
GCEVT [25] 50.6 34.5 520 612 841
GLOA [26] 46.2 39.1 581 824 4426
FOLT [7] 56.9 42.1 - - 800
PID-MOT [37] 50.2 33.0 686 424 3529
AHOR-ReID [21] 56.4 42.5 - - 810
STCMOT [28] 52.0 41.2 667 453 3984
DroneMOT [5] 58.6 43.7 689 397 1112
DFA-MOT [38] 58.3 42.9 518 523 792
MM-Tracker [8] 58.3 44.7 - - -
@@
TABLE II: Tracking performance comparison with state-of-the-art methods on UAVDT test set.
@@
ByteTrack [3] 59.1 41.6 - - 296
UAVMOT [27] 67.3 46.4 624 221 456
OC-SORT [4] 64.9 47.5 - - 288
GCEVT [25] 68.6 47.6 618 363 1801
GLOA [26] 68.9 49.6 626 220 433
FOLT [7] 68.3 48.5 - - 338
STCMOT [28] 69.8 49.2 664 203 665
DroneMOT [5] 69.6 50.1 638 178 129
DFA-MOT [38] 69.3 49.9 684 230 396
MM-Tracker [8] 68.9 51.4 - - -
@@
IV-B2 Results on UAVDT
The tracking scenarios in UAVDT are derived from bird’s-eye views at different outdoor scenes. Table II shows that our AMOT achieves the best tracking performance, with an IDF1 of 74.7% and a MOTA of 55.1%. Specifically, AMOT outperforms STCMOT by +4.9% in IDF1 and surpasses MM-Tracker by +3.7% in MOTA, highlighting its superiority in tracking effectiveness. Moreover, AMOT has the highest MT and the lowest ML, showing its robustness in preserving identity over long-term tracking.
IV-B3 Results on VT-MOT-UAV
VT-MOT-UAV is dedicated to tracking multiple categories, including pedestrians and vehicles, and presents challenges such as varying illumination conditions and cluttered backgrounds. The results are summarized in Table III. Our AMOT exhibits superior performance with an IDF1 of 52.7% and a MOTA of 31.8%, surpassing the previous advanced approaches.
TABLE III: Tracking performance comparison with state-of-the-art methods on VT-MOT-UAV test set.
@@ IV-B3 Results on VT-MOT-UAV
SORT [39] 48.1 28.6 103 340 520
FairMOT [11] 39.9 17.9 94 322 901
ByteTrack [3] 50.2 28.5 129 324 415
UAVMOT [27] 47.9 22.1 175 248 1421
OC-SORT [4] 48.2 28.5 102 342 509
STCMOT [28] 47.4 27.9 153 277 518
@@
IV-C Ablation Studies
IV-C1 Baseline model
The baseline model uses the network architecture of FairMOT [11] combined with a two-stage association strategy [3].
IV-C2 Component Ablation
We conduct a comprehensive evaluation of the proposed components, as presented in Table IV. Integrating either AMC or MTC component into the baseline model individually leads to notable improvements in tracking performance. Compared with the baseline model, the joint application of AMC and MTC improves IDF1 and MOTA by +7.0% and +2.7%, respectively, while reducing identity switches by 2784 on VisDrone2019. On UAVDT, it achieves gains of +2.3% in IDF1 and +0.5% in MOTA, with 614 fewer identity switches. Furthermore, the AMC and MTC introduce only a minimal computational overhead and have a negligible impact on the overall inference speed. Specifically, the baseline model achieves an inference speed of 37.1 FPS, while our AMOT achieves a comparable speed of 36.4 FPS. These results highlight the effectiveness and computational efficiency of the proposed components.
TABLE IV: Ablation of various components.
@@ IV-C2 Component Ablation
TABLE V: Impact of the forward and backward spatial distance in AMC matrix, where “For.” and “Back.” denote forward and backward spatial distances, respectively.
@@
IV-C3 Impact of Bi-directional Spatial Distances
As shown in Table V, we assess the effect of forward and backward spatial distances in the AMC matrix. Both forward and backward spatial distances positively contribute to overall tracking performance, while their joint integration yields further gains. This indicates that modeling bi-directional spatial consistency strengthens the spatio-temporal interplay between motion and appearance cues, thereby improving the discriminative power of the affinity measurement.
Figure 6: Association performance of JDE-based trackers and their variants with our proposed components.
@@ IV-C3 Impact of Bi-directional Spatial Distances
IV-C4 Effect of MTC
Fig. 4 presents a qualitative comparison of tracking results with and without the proposed MTC module. At frame 168, the baseline model fails to associate the object due to missed detections, leading to a track discontinuity. In contrast, the baseline model with MTC module successfully propagates the tracks from frame 167 to frame 168 with consistent identity assignment, highlighting the effectiveness of MTC in preserving track continuity under unreliable detection conditions.
IV-C5 Robustness to Large Displacements
In Fig. 5, we compare the association performance of our AMOT against several advanced data association methods using the same detections under varying frame intervals. As the frame interval increases, inter-frame object displacements become larger, resulting in a significant decline in association accuracy for most methods. In contrast, our AMOT demonstrates strong robustness under these challenging conditions. Although recent competitive methods such as BoT-SORT [40] and Hybrid-SORT [30] exploit both motion and appearance cues for association, they model these cues independently, which limits their ability to handle large object displacements. Instead, our method jointly and adaptively models motion and appearance cues, enabling more consistent and accurate data association in highly dynamic tracking scenarios.
TABLE VI: Results of applying AMC and MTC components to various JDE-based methods.

### v2

@@ Abstract
Multi-object tracking (MOT) aims to track multiple objects while maintaining consistent identities across frames of a given video. In unmanned aerial vehicle (UAV) recorded videos, frequent viewpoint changes and complex UAV-ground relative motion dynamics pose significant challenges, which often lead to unstable affinity measurement and ambiguous association. Existing methods typically model motion and appearance cues separately, overlooking their spatio-temporal interplay and resulting in suboptimal tracking performance. In this work, we propose AMOT, which jointly exploits appearance and motion cues through two key components: an Appearance-Motion Consistency (AMC) matrix and a Motion-aware Track Continuation (MTC) module. Specifically, the AMC matrix computes bi-directional spatial consistency under the guidance of appearance features, enabling more reliable and context-aware identity association. The MTC module complements AMC by reactivating unmatched tracks through appearance-guided predictions that align with Kalman-based predictions, thereby reducing broken trajectories caused by missed detections. Extensive experiments on three UAV benchmarks, including VisDrone2019, UAVDT, and VT-MOT-UAV, demonstrate that our AMOT outperforms current state-of-the-art methods and generalizes well in a plug-and-play and training-free manner.
Codehttps://github.com/ydhcg-BoBo/AMOT
Introduction
Multi-object tracking (MOT) is a fundamental vision task with widespread applications, such as autonomous driving (Zhuang et al. 2024) and unmanned aerial vehicle (UAV) surveillance (Wu et al. 2025). A typical pipeline of MOT is to first detect multiple objects, and then assign each detection to existing tracks through data association, ensuring the continuity of track identities over time. Despite progress, robust data association remains challenging, particularly for videos captured by UAV-mounted cameras.
@@ Introduction
To address the above issues, prevailing methods adopt two main strategies for constructing the cost matrix. (1) Motion-based position prediction: The Kalman Filter (Zhang et al. 2022; Cao et al. 2023) is widely used as a motion model to predict current positions of tracks from their historical states. Then, a motion-based cost matrix is constructed based on the positional proximity between detections and tracks. To overcome the limitations of the Kalman Filter’s linear motion predictions in UAV views, some recent works introduce complementary techniques, such as camera motion compensation (CMC) (Wang et al. 2024a; Song and Lee 2024) and optical flow (Yao et al. 2023, 2025), to improve the accuracy of predicted track positions. (2) Appearance-based instance-level discrimination: Instance-level re-identification (ReID) embedding, being insensitive to position changes, offers significant advantages under substantial camera motion or object displacement. Several works (Li et al. 2024a; Wang et al. 2024a, b; Song and Lee 2024) have constructed appearance-based cost matrices using ReID embedding to facilitate accurate identity assignment. Nevertheless, these two strategies model motion and appearance cues independently to generate separate cost matrices, ignoring the intrinsic relationships between them. Specifically, motion prediction errors caused by sudden object displacements can adversely affect the construction of the motion-based cost matrix, while appearance ambiguities similarly impact the appearance-based cost matrix. Consequently, when these cost matrices produce conflicting association scores, determining a reliable matching decision becomes difficult.
@@
Together with the AMC and MTC modules, we propose a novel multi-object tracker, termed AMOT, which is built on the joint detection and embedding (JDE) architecture (Zhang et al. 2021). AMOT is specifically designed to enhance the robustness and accuracy of identity assignment under challenging UAV-captured videos. Experiments on multiple UAV benchmarks demonstrate that AMOT achieves superior performance in terms of IDF1 and MOTA, as shown in Figure 1. The main contributions are summarized as follows:
@@
Related Works
Motion Modeling
@@
The purpose of motion modeling in MOT is to predict the positions of tracks. Most MOT methods (Stadler and Beyerer 2023; Li et al. 2024b; Cheng et al. 2023) adopt the Kalman Filter for motion modeling, owing to its high computational efficiency for real-time applications. Predicted positions of tracks are compared with current detections via Mahalanobis distance (Du et al. 2023) or Intersection over Union (IoU) (Lv et al. 2024a), which are employed as motion-based cost metrics for data association. Some works (Cao et al. 2023; Maggiolino et al. 2023) enhance the Kalman Filter by adopting an observation-centric strategy to refine track estimation, yielding better performance under non-linear motion patterns. Despite their effectiveness, Kalman-based methods exhibit limited capability in scenarios involving substantial camera motion and object displacement. Meanwhile, learning-based motion models (Shuai et al. 2021; Qin et al. 2023; Song et al. 2025) have recently gained attention. These models leverage data-driven architectures to learn motion patterns. However, such methods often incur high computational costs and are unsuitable for real-time tracking. In contrast, we introduce a training-free motion modeling strategy that balances robustness and efficiency, specifically tailored for UAV-captured videos.
Appearance Modeling
@@
Appearance modeling aims to extract discriminative appearance features to re-identify objects. The tracking-by-detection paradigm (Jin et al. 2024; Huang et al. 2024) first detects objects using an off-the-shelf detector, and then employs a ReID network to extract identity embedding for each object. Although this pipeline delivers impressive performance, it suffers from high computational cost due to separate detection and extraction stages. In contrast, the joint detection and embedding (JDE) paradigm (Liu et al. 2023; Meng et al. 2023) integrates object detection and ReID feature extraction into a unified framework, enabling simultaneous object localization and embedding extraction. Furthermore, several JDE-based trackers incorporate global attention (Wu et al. 2023; Shi et al. 2023) and temporal cues (Liu et al. 2022; Ma et al. 2024) to enhance the discriminability of instance-level ReID embedding. The aforementioned methods measure instance-level appearance similarity by computing the cosine distance between ReID embeddings. Differently, we reformulate the appearance similarity measurement as a dense response between the instance embedding and the global ReID feature map. This enables the joint modeling of visual similarity and spatial coherence.
@@
Figure 2: Tracking pipeline of AMOT. Specifically, we introduce the appearance-motion consistency (AMC) matrix $\mathbf{C}_{AMC}$ that integrates it with the appearance similarity matrix $\mathbf{C}_{App}$ and the Intersection-over-Union (IOU) matrix $\mathbf{C}_{IOU}$ to robustly associate high-confidence detections $\mathcal{D}_{h}$ with tracks $\mathcal{T}$ at frame t1t-1. Then, the unmatched tracks $\mathcal{T}_{un}$ are associated with low-confidence detections $\mathcal{D}_{l}$ in the second stage. The remaining unmatched tracks $\mathcal{T}_{second\_un}$ are further potentially reactivated through our proposed motion-aware track continuation (MTC) module. KF means the Kalman Filter.
Data Association
@@
Data association typically involves constructing a cost matrix that quantifies the affinity between current detections and existing tracks based on motion and appearance information. To improve association accuracy, motion and appearance cues are often modeled separately and then integrated into a unified cost matrix (Liang et al. 2022; Yang et al. 2024). Once the cost matrix is constructed, data association is formulated as an assignment problem and solved using the Hungarian algorithm (Yi et al. 2024; Qin et al. 2024). Nevertheless, existing approaches commonly fail to consider the inherent interplay between motion and appearance cues. This independent modeling can cause instability in affinity measurement, arising from prediction errors or appearance ambiguities, which ultimately leads to conflicting data associations. To this end, we propose the AMC matrix, which jointly enforces consistency in appearance, motion, and temporal domains, enabling reliable identity assignment.
Methodology
Preliminaries
@@
where $\mathcal{O}_{det}$ denotes the set of center coordinates for detections, and $N$ is the total number of detections. For each detection, the width and height are obtained from the regression map as $w_{i},h_{i}]=\textbf{B}^{t}_{(x_{i},y_{i})}$, and the confidence score is given by $c_{i}=\mathbf{H}^{t}_{(x_{i},y_{i})}$. Meanwhile, the ReID embedding $\mathbf{e}_{i}$ for each detection is extracted from $\textbf{E}^{t}$ at the corresponding center coordinate, expressed as $\mathbf{e}_{i}=\mathbf{E}^{t}_{(x_{i},y_{i})}\in\mathbb{R}^{D}$.
@@
In our AMOT, as shown in Figure 2, we construct an appearance-motion consistency (AMC) matrix to enable robust detection-to-track association. Additionally, we propose a motion-aware track continuation (MTC) module designed to reactivate unmatched tracks without relying on explicit detection-to-track matching.
Appearance-Motion Consistency Matrix
@@
where $\mathbf{e}_{j}$ denotes the ReID embedding of the $j$-th track, and $\text{sim}(\cdot)$ is the cosine similarity function. The response map $\mathbf{A}_{trk}^{\scriptscriptstyle(j)}\in\mathbb{R}^{H\times W}$ highlights regions that are most semantically correlated with the track’s ReID embedding. Then, the spatial location corresponding to the maximum response on $\mathbf{A}_{trk}^{\scriptscriptstyle(j)}$ is defined as the center of the $j$-th track in the current frame. Accordingly, the set of predicted center coordinates $\mathcal{Q}_{{trk}}$ for tracks in the current frame is formulated as:
@@
Subsequently, as shown in Figure 3, we quantify the appearance-guided spatio-temporal correspondence by measuring the forward and backward spatial distances, which can be defined as:
@@
Motion-aware Track Continuation
@@
Tracking Pipeline
The tracking pipeline of AMOT is illustrated in Figure 2. The Kalman Filter is used to predict the current positions of tracks. We then adopt a two-stage matching strategy similar to BYTE (Zhang et al. 2022), where detections are divided into high-confidence sets $\mathcal{D}_{h}$ and low-confidence sets $\mathcal{D}_{l}$.
Specifically, in the first stage, we construct three pairwise cost matrices between $\mathcal{D}_{h}$ and the tracks $\mathcal{T}$, including the appearance similarity matrix $\mathbf{C}_{App}$ (Zhang et al. 2021), the Intersection-over-Union (IOU) matrix $\mathbf{C}_{IOU}$, and the proposed AMC matrix $\mathbf{C}_{AMC}$. These matrices are integrated into a unified cost matrix $\mathbf{C}_{uni}$ as follows:
@@
Experiments
Settings
Datasets and Evaluation Metrics.
We evaluated our method on various MOT datasets from UAV perspectives, including VisDrone2019 (Du et al. 2019), UAVDT (Du et al. 2018), and VT-MOT-UAV (Zhu et al. 2025).
We adopt the widely-used CLEAR metrics (Liu et al. 2022), including MOTA, IDF1, the number of mostly tracked (MT) objects, and mostly lost (ML) objects, to comprehensively evaluate the tracking performance. MOTA emphasizes the detection quality and is computed based on false positives (FP), false negatives (FN), and identity switches (IDs). IDF1 measures the tracker’s ability to maintain consistent object identities over time, reflecting the accuracy of identity association across frames. Additionally, we report FPS to assess the inference speed of the tracker.
Implementation Details.
@@
Benchmark Evaluation
We present the tracking results for multiple UAV datasets. $\uparrow$$\downarrow$ indicate that higher/lower is better, respectively. The best scores for each metric are marked in bold, and the second-best scores are marked in underline.
[…]
ByteTrack (Zhang et al. 2022) 37.0 35.7 - - 2168
UAVMOT (Liu et al. 2022) 51.0 36.1 520 574 2775
OC-SORT (Cao et al. 2023) 50.4 39.6 - - 986
GCEVT (Wu et al. 2023) 50.6 34.5 520 612 841
GLOA (Shi et al. 2023) 46.2 39.1 581 824 4426
FOLT (Yao et al. 2023) 56.9 42.1 - - 800
PID-MOT (Lv et al. 2024b) 50.2 33.0 686 424 3529
AHOR-ReID (Jin et al. 2024) 56.4 42.5 - - 810
STCMOT (Ma et al. 2024) 52.0 41.2 667 453 3984
DroneMOT (Wang et al. 2024a) 58.6 43.7 689 397 1112
DFA-MOT (Zheng et al. 2025) 58.3 42.9 518 523 792
MM-Tracker (Yao et al. 2025) 58.3 44.7 - - -
@@
Table 1: Tracking performance comparison with state-of-the-art methods on VisDrone2019 test set.
@@
ByteTrack (Zhang et al. 2022) 59.1 41.6 - - 296
UAVMOT (Liu et al. 2022) 67.3 46.4 624 221 456
OC-SORT (Cao et al. 2023) 64.9 47.5 - - 288
GCEVT (Wu et al. 2023) 68.6 47.6 618 363 1801
GLOA (Shi et al. 2023) 68.9 49.6 626 220 433
FOLT (Yao et al. 2023) 68.3 48.5 - - 338
STCMOT (Ma et al. 2024) 69.8 49.2 664 203 665
DroneMOT (Wang et al. 2024a) 69.6 50.1 638 178 129
DFA-MOT (Zheng et al. 2025) 69.3 49.9 684 230 396
MM-Tracker (Yao et al. 2025) 68.9 51.4 - - -
@@
Table 2: Tracking performance comparison with state-of-the-art methods on UAVDT test set.
Results on VisDrone2019.
VisDrone2019 serves as a fundamental benchmark for multi-object tracking in dynamic UAV-captured videos, which involve scale variations and long-range object movements. As shown in Table 1, our AMOT exhibits the highest IDF1 of 61.4% and MOTA of 46.0%, outperforming current state-of-the-art methods. Concretely, compared with DroneMOT and MM-Tracker, AMOT achieves absolute improvements in both IDF1 and MOTA, demonstrating superior association accuracy and enhanced overall tracking performance.
Results on UAVDT.
The tracking scenarios in UAVDT are derived from bird’s-eye views at different outdoor scenes. Table 2 shows that our AMOT achieves the best tracking performance, with an IDF1 of 74.7% and a MOTA of 55.1%. Specifically, AMOT outperforms STCMOT by +4.9% in IDF1 and surpasses MM-Tracker by +3.7% in MOTA, highlighting its superiority in tracking effectiveness. Moreover, AMOT has the highest MT and the lowest ML, showing its robustness in preserving identity over long-term tracking.
Results on VT-MOT-UAV.
VT-MOT-UAV is dedicated to tracking multiple categories, including pedestrians and vehicles, and presents challenges such as varying illumination conditions and cluttered backgrounds. The results are summarized in Table 3. Our AMOT exhibits superior performance with an IDF1 of 52.7% and a MOTA of 31.8%, surpassing the previous advanced approaches.
@@
SORT (Bewley et al. 2016) 48.1 28.6 103 340 520
FairMOT (Zhang et al. 2021) 39.9 17.9 94 322 901
ByteTrack (Zhang et al. 2022) 50.2 28.5 129 324 415
UAVMOT (Liu et al. 2022) 47.9 22.1 175 248 1421
OC-SORT (Cao et al. 2023) 48.2 28.5 102 342 509
STCMOT (Ma et al. 2024) 47.4 27.9 153 277 518
@@
Table 3: Tracking performance comparison with state-of-the-art methods on VT-MOT-UAV test set.
@@
Table 4: Ablation of various components.
Ablation Studies
Baseline model.
The baseline model uses the network architecture of FairMOT (Zhang et al. 2021) combined with a two-stage association strategy (Zhang et al. 2022).
@@
Table 5: Impact of the forward and backward spatial distance in AMC matrix, where “For.” and “Back.” denote forward and backward spatial distances, respectively.
Component Ablation.
We conduct a comprehensive evaluation of the proposed components, as presented in Table 4. Integrating either AMC or MTC component into the baseline model individually leads to notable improvements in tracking performance. Compared with the baseline model, the joint application of AMC and MTC improves IDF1 and MOTA by +7.0% and +2.7%, respectively, while reducing identity switches by 2784 on VisDrone2019. On UAVDT, it achieves gains of +2.3% in IDF1 and +0.5% in MOTA, with 614 fewer identity switches. Furthermore, the AMC and MTC introduce only a minimal computational overhead and have a negligible impact on the overall inference speed. Specifically, the baseline model achieves an inference speed of 37.1 FPS, while our AMOT achieves a comparable speed of 36.4 FPS. These results highlight the effectiveness and computational efficiency of the proposed components.
Impact of Bi-directional Spatial Distances.
As shown in Table 5, we assess the effect of forward and backward spatial distances in the AMC matrix. Both forward and backward spatial distances positively contribute to overall tracking performance, while their joint integration yields further gains. This indicates that modeling bi-directional spatial consistency strengthens the spatio-temporal interplay between motion and appearance cues, thereby improving the discriminative power of the affinity measurement.
@@
Effect of MTC.
Figure 4 presents a qualitative comparison of tracking results with and without the proposed MTC module. At frame 168, the baseline model fails to associate the object due to missed detections, leading to a track discontinuity. In contrast, the baseline model with MTC module successfully propagates the tracks from frame 167 to frame 168 with consistent identity assignment, highlighting the effectiveness of MTC in preserving track continuity under unreliable detection conditions.
Robustness to Large Displacements.
In Figure 5, we compare the association performance of our AMOT against several advanced data association methods using the same detections under varying frame intervals. As the frame interval increases, inter-frame object displacements become larger, resulting in a significant decline in association accuracy for most methods. In contrast, our AMOT demonstrates strong robustness under these challenging conditions. Although recent competitive methods such as BoT-SORT (Aharon et al. 2022) and Hybrid-SORT (Yang et al. 2024) exploit both motion and appearance cues for association, they model these cues independently, which limits their ability to handle large object displacements. Instead, our method jointly and adaptively models motion and appearance cues, enabling more consistent and accurate data association in highly dynamic tracking scenarios.